# From Logits to Hierarchies: Hierarchical Clustering made Simple

### v1

@@ Abstract
The structure of many real-world datasets is intrinsically hierarchical, making the modeling of such hierarchies a critical objective in both unsupervised and supervised machine learning. Recently, novel approaches for hierarchical clustering with deep architectures have been proposed. In this work, we take a critical perspective on this line of research and demonstrate that many approaches exhibit major limitations when applied to realistic datasets, partly due to their high computational complexity. In particular, we show that a lightweight procedure implemented on top of pre-trained non-hierarchical clustering models outperforms models designed specifically for hierarchical clustering. Our proposed approach is computationally efficient and applicable to any pre-trained clustering model that outputs logits, without requiring any fine-tuning. To highlight the generality of our findings, we illustrate how our method can also be applied in a supervised setup, recovering meaningful hierarchies from a pre-trained ImageNet classifier.
@@ 1 Introduction
Modeling hierarchical structures in the data is a long-standing goal in machine learning research (Bengio et al., 2014; Jordan & Mitchell, 2015). In many real-world scenarios, data is inherently organized in hierarchies, such as phylogenetic trees (Linnæus, 1758; Sneath & Sokal, 1962; Penny, 2004), tumor subclasses (Sørlie et al., 2001) and social networks Ravasz & Barabási (2003); Crockett et al. (2017). In unsupervised learning, hierarchical clustering can provide more accurate insights than flat (i.e. non-hierarchical) clustering methods by introducing multiple levels of granularity and alleviating the need for a fixed number of clusters specified a priori (Bertsimas et al., 2021; Chami et al., 2020). This aids scientific understanding and interpretability by providing a more informative representation (Lipton, 2018; Marcinkevičs & Vogt, 2020). The benefits of modeling a hierarchy in the data extend to supervised scenarios. For example, interpretable methods based on decision trees (Breiman, 2001; Tanno et al., 2019) hierarchically partition the data so that points in each split are linearly separable into classes. More recent work leverages hierarchies in the data to improve supervised methods (Bertinetto et al., 2020; Goren et al., 2024; Karthik et al., 2021) or for self-supervision (Long & van Noord, 2023).
@@
Among classic algorithms for hierarchical clustering, agglomerative methods have been the most widely adopted. These methods compute pairwise distances between data points, often in a lower-dimensional representation space. Starting from the instance level, a hierarchy is then built based on the pairwise distances by recursive agglomeration of similar points or clusters together in a bottom-up fashion (Murtagh & Contreras, 2011). More recently, a revived interest in hierarchical clustering has sparked novel approaches using deep architectures (Mautz et al., 2020; Goyal et al., 2017; Shin et al., 2019; Vikram et al., 2019; Manduchi et al., 2023). However, despite their promising methodological contributions, recent approaches require specialized architectures and complex training schemes. Consequently, they cannot be applied to large-scale datasets due to their expensive computational requirements. Moreover, we find that they often exhibit a lower performance at the leaf level compared to non-hierarchical models.
@@
In this work, we take a critical perspective on recent research on hierarchical clustering and offer a simple alternative. Instead of designing specialized hierarchical clustering models, we develop a lightweight method for hierarchical clustering given a pre-trained flat model. In particular, we show that a lightweight algorithm implemented on top of (non-hierarchical) pre-trained models markedly outperforms specialized models for hierarchical clustering. Notably, our algorithm, which we name Logits to Hierarchies (L2H), only uses logits and requires no fine-tuning of the pre-trained model. Hence, it generally applies to black-box models even without access to internal representations (e.g., API calls to proprietary models) and bypasses the costly computation of pairwise distances between data points. Moreover, it also applies to supervised models, for which the inferred hierarchy of classes can aid model interpretability, e.g., for discovering potential biases such as spurious correlations between classes.
@@
  • In Section 3, we propose a simple algorithm for hierarchical clustering that transforms the logits from a pre-trained model into a hierarchical structure of classes. Our method markedly outperforms specialized hierarchical models and has low computational requirements. With logits as input, it computes a hierarchical clustering on ImageNet-sized datasets in a few minutes on a single CPU core.
  • In Section 4.1, our experiments reveal significant limitations of recently proposed methods for hierarchical clustering, highlighting their weaknesses on large-scale datasets and subpar performance at the leaf level compared to non-hierarchical approaches.
  • In Section 4.2, we provide a case study on ImageNet to demonstrate how our method applies to supervised models, showing how the inferred hierarchy of classes recovers parts of the WordNet hierarchy and helps discover potential biases of the pre-trained model and ambiguities in existing categorizations.
@@ 2 Related work
Hierarchical clustering aims to learn clusters of datapoints that are organized in a hierarchical structure. The methods used can be broadly categorized into agglomerative and divisive approaches (Nielsen, 2016). The former tackles the problem with a bottom-up approach and iteratively agglomerates clusters into larger ones until a full hierarchy is built in the form of a dendrogram, starting with each datapoint being a separate cluster (Murtagh & Contreras, 2011). The similarity of datapoints is measured according to a distance function, which for high-dimensional data is often defined on a lower-dimensional representation space. Multiple linkage methods have been proposed to compute the distance between clusters of datapoints formed at a given step of the algorithm (Sneath, 1957; Ward, 1963). As examples, single, average, and complete linkage characterize the distance between two clusters as the minimum, average, and maximum distance between their datapoints, respectively. Since these algorithms can be costly, particularly in high-dimensional spaces, approximate versions have been developed for faster computation (Abboud et al., 2019; Cochez & Mou, 2015). Notably, linkage methods are still widely applied in many domains, for instance, in medical research (Nguyen et al., 2024; Senevirathna et al., 2023; Resende et al., 2023).
@@
On the other hand, divisive algorithms start with all objects belonging to the same cluster and recursively split them into subclusters. While early approaches are mostly based on heuristics, Dasgupta. (2016) proposed the Dasgupta cost: an objective function for evaluating a hierarchical clustering, with a divisive approach to provide an approximately optimal solution. HypHC introduces a continuous relaxation of Dasgupta’s discrete optimization problem with provable guarantees via hyperbolic embeddings that better reflect the geometry of trees compared to Euclidean representations Chami et al. (2020); Liu et al. (2019). More recently, research has focused on developing deep learning approaches for hierarchical clustering (Mautz et al., 2020; Goyal et al., 2017; Shin et al., 2019; Vikram et al., 2019; Manduchi et al., 2023). Among these, DeepECT learns a hierarchical clustering on top of the embedding space of a jointly optimized autoencoder (Mautz et al., 2020). Notably, TreeVAE not only learns a hierarchical clustering in the latent space but also provides a generative model that adheres to the learnt hierarchy, thereby enabling sample generation in a structured manner (Manduchi et al., 2023). However, all these approaches have mostly been tested on simple datasets, far from realistic settings. As shown in our experiments, they present relevant limitations when deployed on more challenging datasets, mainly due to their high computational complexity.
@@
Finally, the benefits of modeling a hierarchy in the data are not restricted to the unsupervised setup. In particular, a recent line of research focuses on leveraging a tree structure in the classes to assess and reduce the severity of misclassification of supervised models (Karthik et al., 2021). This can lead to safer models in cost-sensitive classification scenarios (Bertinetto et al., 2020) and allow a classifier to predict at different levels of the hierarchy depending on the required confidence (Goren et al., 2024). The visualization of hierarchies also provides global explanations of a model’s functionality, thereby improving a user’s understanding of the model behavior and fostering trust (Chakraborty et al., 2017; Lipton, 2018).
@@ 3 Method
In this work, we take a critical perspective on a recent line of research on hierarchical clustering: as an alternative to designing ad-hoc complex approaches, we focus on adapting pre-trained flat models to output a hierarchy with minimal overhead. To this end, we introduce a lightweight algorithm to leverage the information contained in the logits of a pre-trained flat clustering model to output a hierarchy of clusters. In the following, we describe the proposed procedure and also provide a graphical illustration as well as detailed pseudocode.
Let $\mathcal{D}=\{\bm{x}_{1},\dots,\bm{x}_{N}\}$ be a dataset consisting of $N$ data points and $f_{\theta}$ be a non-hierarchical model trained to partition $\mathcal{D}$ into $K$ clusters. We assume that $f_{\theta}$ outputs unnormalized logits, i.e. the cluster assignment $k^{*}$ for a datapoint $\bm{x}$ is determined by computing $k^{*}=\operatorname*{arg\,max}_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))$. We define two functions
$h_{\theta}(\bm{x})=\operatorname*{arg\,max}_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))\qquad\qquad g_{\theta}(\bm{x})=\max_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))$
@@
A key idea behind our method is a simple yet effective way to determine the relatedness of clusters, or groups of clusters, by iteratively grouping them together to construct a hierarchy. Intuitively, to assess which group of clusters $G^{\prime}$ is most related to a given group $G$, we propose the following strategy. For datapoints assigned to clusters in $G$, we determine which group $G^{\prime}$ would have the majority of these datapoints reassigned to if clusters in $G$ were not available. [^1] This passage is primarily for intuition and not strictly accurate. To be precise, we look not at reassignments but at predicted probabilities of reassignments (see Equation 1).compute cluster assignments and corresponding predicted probabilities, restricting only to a subset of the total set of clusters.
@@
[…]
given a $K$-dimensional vector $\bm{v}$ and a set $G\subset\{1,\dots,K\}$. This function restricts the softmax operation to the elements of $\bm{v}$ at indexes in $\{1,\dots,K\}\setminus G$. Next, we define functions
$h^{m}_{\theta}(\bm{x};G)=\operatorname*{arg\,max}_{k\in\{1,\dots,K\}}\text{m\_softmax}_{k}(f_{\theta}(\bm{x});G)\qquad g^{m}_{\theta}(\bm{x};G)=\max_{k\in\{1,\dots,K\}}\text{m\_softmax}_{k}(f_{\theta}(\bm{x});G)$
Note that the $h^{m}_{\theta}$ and $g^{m}_{\theta}$ functions correspond to $h_{\theta}$ and $g_{\theta}$, except restricting the choice of viable clusters to $\{1,\dots,K\}\setminus G$. In particular, the $h^{m}_{\theta}$ function computes the cluster assignment for a datapoint $\bm{x}$ restricting to clusters in $\{1,\dots,K\}\setminus G$, and $g^{m}_{\theta}$ the corresponding predicted probability. Lastly, we define
$\mathcal{D}^{c}:=\{\bm{x}\in\mathcal{D}\mid h_{\theta}(\bm{x})=c\}$
i.e. the subset of datapoints assigned to a given cluster $c\in\{1,\dots,K\}$. Similarly, we denote as $\mathcal{D}^{G}=\cup_{c\in G}D^{c}$ the subset of datapoints assigned to a group of clusters $G\subset\{1,\dots,K\}$.
Algorithm 1 Logits to Hierarchies (L2H)
  Input: Pre-trained model $f_{\theta}$ with parameters $\theta$, Dataset $\mathcal{D}$.
@@
  Initialize groups $\mathcal{G}=\{G_{1},\dots,G_{K}\}$ where $G_{k}=\{k\}$     # Groups initialized as single clusters
  H = []     # Hierarchy initialized as empty list
  for step $s$ from $1$ to $K-1$ do
@@
     Take $G^{\star}\in\text{argmin}_{G\in\mathcal{G}}s(G)$     # Select group with lowest score for merging
@@
        Compute $rp(c):=\sum_{\begin{subarray}{c}\bm{x}\in\mathcal{D}^{G^{*}}\\ h_{\theta}^{m}(\bm{x};G^{*})=c\end{subarray}}g_{\theta}^{m}(\bm{x};G^{*})$     # Compute total predicted probability per cluster
@@
     Take $G^{\dagger}\in\text{argmax}_{G\in\mathcal{G}\setminus\{G^{*}\}}\frac{1}{|G|}\sum_{c\in G}rp(c)$     # Select $G^{\dagger}$, most related group to $G^{*}$, for merging
     Update $\mathcal{G}$ by merging groups $G^{*}$ and $G^{\dagger}$     # Update groups
     Update H by adding that $G^{*}$ and $G^{\dagger}$ are merged at step $s$     # Update hierarchy
@@
We describe our proposed method in Algorithm 1. At the start of the procedure, $K$ groups are initialized as single clusters. At each iteration, two groups are merged into a single group, constructing a tree of clusters up to the root in $K-1$ iterations. Each iteration can be split in two stages. In the first stage, a score is computed for each group. To compute the score for a given group $G$ we aggregate the predicted probabilities for the datapoints assigned to clusters contained in $G$ as
@@
where $\Lambda$ is a chosen aggregation function (e.g. sum function). Then, the lowest scored group $G^{*}$ is selected for merging at this iteration, which concludes the first stage.
In the second stage, we search for the group $G^{\dagger}$ that is most related to $G^{*}$ to perform the merging. To do so, as mentioned above, we look at the subset of datapoints assigned to clusters in $G^{*}$: for these datapoints we recompute cluster assignments and predicted probabilities, this time restricting to clusters not contained in $G^{*}$. More formally, the total reassigned predicted probability to each cluster not contained in $G^{*}$ is computed as
$rp(c):=\sum_{\begin{subarray}{c}\bm{x}\in\mathcal{D}^{G^{*}}\\ h_{\theta}^{m}(\bm{x};G^{*})=c\end{subarray}}g_{\theta}^{m}(\bm{x};G^{*})\qquad\forall c\in\{1,..,K\}\setminus G^{*}$ (1)
Note that this quantity can be interpreted as a measure of relatedness between each cluster $c\in\{1,..,K\}\setminus G^{*}$, and the group of clusters $G^{*}$. The most related group to $G^{*}$ is finally selected as $G^{\dagger}\in\text{argmax}_{G\in\mathcal{G}\setminus\{G^{*}\}}\frac{1}{|G|}\sum_{c\in G}rp(c)$, i.e. by averaging the total reassigned predicted probability across clusters in each group and selecting the group with the highest average. Given that cluster assignments and corresponding predicted probabilities can be computed via simple operations on the logits, the whole procedure can be executed with as input only the logits for the dataset $\mathcal{D}$ outputted from a pre-trained model $f_{\theta}$. To show this, we report an example Python implementation in Appendix A. Additionally Figure 1 provides an illustration to exemplify the proposed grouping strategy.

### v2

@@ Abstract
The hierarchical structure inherent in many real-world datasets makes the modeling of such hierarchies a crucial objective in both unsupervised and supervised machine learning. While recent advancements have introduced deep architectures specifically designed for hierarchical clustering, we adopt a critical perspective on this line of research. Our findings reveal that these methods face significant limitations in scalability and performance when applied to realistic datasets. Given these findings, we present an alternative approach and introduce a lightweight method that builds on pre-trained non-hierarchical clustering models. Remarkably, our approach outperforms specialized deep models for hierarchical clustering, and it is broadly applicable to any pre-trained clustering model that outputs logits, without requiring any fine-tuning. To highlight the generality of our approach, we extend its application to a supervised setting, demonstrating its ability to recover meaningful hierarchies from a pre-trained ImageNet classifier. Our results establish a practical and effective alternative to existing deep hierarchical clustering methods, with significant advantages in efficiency, scalability and performance.
Machine Learning, ICML
@@ 1 Introduction
Modeling hierarchical structures in the data is a long-standing goal in machine learning research (hierclustmotivation1; hierclustmotivation2). In many real-world scenarios, data is inherently organized in hierarchies, such as phylogenetic trees (linnaeus1758systema; hierclust_phylogeny_1; hierclust_phylogeny_2), tumor subclasses (hierclust_sequencing_1) and social networks (ravasz2003hierarchical; crockett2017cluster). In unsupervised learning, hierarchical clustering can provide more accurate insights than flat (i.e., non-hierarchical) clustering methods by introducing multiple levels of granularity and alleviating the need for a fixed number of clusters specified a priori (bertsimas2021interpretable; chami2020trees). This aids scientific understanding and interpretability by providing a more informative representation (lipton2018mythos; marcinkevivcs2020interpretability). The benefits of modeling hierarchies in the data extend to supervised scenarios. For example, interpretable methods based on decision trees (randomforest_paper; tanno2019adaptiveneuraltrees) hierarchically partition the data so that points in each split are linearly separable into classes. More recent work leverages hierarchies in the data to improve supervised methods (bertinetto2020makingbettermistakesleveraging; goren2024hierarchicalselectiveclassification; karthik2021costlikelihoodmanipulationtest) or for self-supervision (long2023).
@@
Among classic algorithms for hierarchical clustering, agglomerative methods have been the most widely adopted. These methods compute pairwise distances between data points, often in a lower-dimensional representation space. Starting from the instance level, a hierarchy is then built based on the pairwise distances by recursive agglomeration of similar points or clusters together in a bottom-up fashion (murtagh2011algorithms). More recently, a revived interest in hierarchical clustering has sparked novel, sophisticated approaches using deep architectures (mautz2020deep; goyal2017nonparametric; shin2019hierarchical; vikram2019theloracs; manduchi2023tree). These approaches require specialized architectures and complex training schemes. In this work, we uncover two key limitations of these approaches. First, we find that these methods struggle to handle large-scale datasets, failing to deliver satisfactory performance. This is in part due to their high computational demands and in part due to the difficulty in modeling a large number of clusters. Second, we observe a notable gap in performance at the leaf level compared to non-hierarchical models. This is particularly problematic, as the advantage of introducing a hierarchy should not come at the expense of the quality of clustering at the leaf-level granularity. These limitations underscore the need for more scalable and effective approaches for hierarchical clustering.
@@
Given these findings, we take a critical perspective on recent research on deep hierarchical clustering and offer a straightforward yet effective alternative. Rather than designing specialized hierarchical clustering models, we develop a lightweight method to perform hierarchical clustering given a pre-trained flat model. In particular, we show that a lightweight algorithm implemented on top of (non-hierarchical) pre-trained models produces hierarchical clustering results that markedly outperform state-of-the-art dedicated models. Notably, our algorithm, which we name Logits to Hierarchies (L2H), only uses logits and requires no fine-tuning of the pre-trained model. Hence, it generally applies to black-box models even without access to internal representations (e.g., API calls to proprietary models) and bypasses the costly computation of pairwise distances between data points. Moreover, it also applies to supervised models, for which the inferred hierarchy of classes can aid model interpretability, e.g., for discovering potential biases such as spurious correlations between classes.
@@
  • We reveal significant limitations of recently proposed deep specialized methods for hierarchical clustering, highlighting their weaknesses on large-scale datasets and their subpar performance at the leaf level compared to non-hierarchical approaches.
  • As an alternative approach, we propose a straightforward algorithm for hierarchical clustering that from the logits of a pre-trained flat model derives a hierarchical structure of clusters. Our method markedly outperforms specialized hierarchical models and has low computational requirements. With logits as its input, it computes a hierarchical clustering on ImageNet-sized datasets in under a minute on a CPU.
  • To demonstrate how our method also applies to supervised models, we provide a case study on ImageNet, showing how the inferred hierarchy of classes recovers parts of the WordNet hierarchy, and helps to discover potential spurious correlations in the pre-trained model or biases in existing categorizations.
@@ 2 Related Work
Hierarchical clustering aims to learn clusters of data points that are organized in a hierarchical structure. The methods used can be broadly categorized into agglomerative and divisive approaches (nielsen2016hierarchical). The former tackle the problem with a bottom-up approach and iteratively agglomerate clusters into larger ones until a full hierarchy is built in the form of a dendrogram, starting with each datapoint being a separate cluster (murtagh2011algorithms). The similarity of data points is measured according to a distance function, which for high-dimensional data is often defined on a lower-dimensional representation space. Multiple linkage methods have been proposed to compute the distance between clusters of data points formed at a given step of the algorithm (sneath1957application; Ward1963HierarchicalGT). As examples, single, average, and complete linkage characterize the distance between two clusters as the minimum, average, and maximum distance between their data points, respectively. Since these algorithms can be costly, particularly in high-dimensional spaces, approximate versions have been developed for faster computation (ApproxAggHC1; ApproxAggHC2). Notably, linkage methods are still widely applied in many domains as, for instance, in medical research (Nguyen:2024aa; Senevirathna:2023aa; Resende:2023aa). Recent work also adopts them to assess how well the representations from pre-trained encoders generalize to cluster unseen datasets (lowe2024empiricalstudyclusteringunseen).
@@
On the other hand, divisive algorithms start with all objects belonging to the same cluster and recursively split them into subclusters. While early approaches are mostly based on heuristics, dasgupta2016acost proposed the Dasgupta cost: an objective function for evaluating a hierarchical clustering, with a divisive approach to provide an approximately optimal solution. HypHC introduces a continuous relaxation of Dasgupta’s discrete optimization problem with provable guarantees via hyperbolic embeddings that better reflect the geometry of trees compared to Euclidean representations (chami2020trees; liu2019hyperbolic). More recently, research has focused on developing deep learning approaches for hierarchical clustering with specialized architectures (mautz2020deep; goyal2017nonparametric; shin2019hierarchical; vikram2019theloracs; manduchi2023tree). Among these, DeepECT learns a hierarchical clustering on top of the embedding space of a jointly optimized autoencoder (mautz2020deep). TreeVAE (manduchi2023tree), on the other hand, learns a hierarchical clustering in the latent space of a variational autoencoder and provides a generative model that adheres to the learned hierarchy, thereby enabling sample generation in a structured manner (manduchi2023tree). However, these approaches have mostly been tested on simple datasets, far from realistic settings. In our experiments in Section 4.1, we show that they present important limitations when deployed on more challenging datasets. We find these limitations to be linked to their high computational demands and their difficulty in modeling hierarchies that consist of a large number of leaf clusters.
@@
Finally, the benefits of modeling a hierarchy in the data are not restricted to the unsupervised setup (hypimageembeddings2020; linderman2022finegraininferenceoutofdistributiondata; sinha2024learning). In particular, recent research focuses on leveraging a tree structure in the classes to assess and reduce the severity of misclassification of supervised models (karthik2021costlikelihoodmanipulationtest). This can lead to safer models in cost-sensitive scenarios (bertinetto2020makingbettermistakesleveraging) and allow a classifier to predict at different levels of the hierarchy depending on the required confidence (goren2024hierarchicalselectiveclassification). The visualization of hierarchies also provides global explanations of a model’s functionality, thereby improving a user’s understanding of the model behavior and fostering trust (chakraborty2017interpretability; lipton2018mythos).
@@ 3 Method
In this work, we take a critical perspective on a recent line of research on hierarchical clustering. After uncovering important limitations of recent hierarchical clustering approaches with deep specialized architectures, we propose an alternative strategy. Namely, rather than designing ad-hoc complex methods, we focus on adapting pre-trained flat models to output a hierarchy with minimal overhead. To this end, we introduce a lightweight algorithm to leverage the information contained in the logits of a pre-trained flat clustering model to output a hierarchy of clusters. In the following, we describe the proposed procedure and also provide a graphical illustration as well as detailed pseudocode.
Let $\mathcal{D}=\{\bm{x}_{1},\dots,\bm{x}_{N}\}$ be a dataset consisting of $N$ data points and $f_{\theta}$ be a non-hierarchical model trained to partition $\mathcal{D}$ into $K$ clusters. We assume that $f_{\theta}$ outputs unnormalized logits, from which the cluster assignment $k^{*}$ for a datapoint $\bm{x}$ is determined by taking the softmax
$k^{*}=\operatorname*{argmax}_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))\;.$
We define two functions
$h_{\theta}(\bm{x})=\operatorname*{argmax}_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))\;,$
$g_{\theta}(\bm{x})=\max_{k\in\{1,\dots,K\}}\text{softmax}_{k}(f_{\theta}(\bm{x}))\;,$
@@
A key idea behind our method is a simple yet effective way to determine the relatedness of clusters, or groups of clusters, while iteratively grouping them together to construct a hierarchy. Intuitively, to assess which group of clusters $G^{\prime}$ is most related to a given group $G$, we propose the following strategy: for data points assigned to clusters in $G$, we determine which group $G^{\prime}$ would have the majority of these data points reassigned to, if clusters in $G$ were not available. [^1] This passage is primarily for intuition and not strictly accurate. To be precise, we do not look at reassignments but rather at predicted probabilities of reassignments (see Equation 2). Formally, we define the following functions to compute cluster assignments and corresponding predicted probabilities, restricting only to a subset of the total set of clusters.
@@
$\text{m\_softmax}_{k}(\bm{v};G)=\begin{cases}\frac{\text{exp}({v_{i}})}{\sum_{j\in\{1,\dots,K\}\setminus G}\text{exp}({v_{j}})}&\text{if }i\notin G\\ 0&\text{if }i\in G\end{cases}$
given a $K$-dimensional vector $\bm{v}$ and a set $G\subset\{1,\dots,K\}$. This function restricts the softmax operation to the elements of $\bm{v}$ at indexes in $\{1,\dots,K\}\setminus G$. Next, we define
$h^{m}_{\theta}(\bm{x};G)=\operatorname*{argmax}_{k\in\{1,\dots,K\}}\text{m\_softmax}_{k}(f_{\theta}(\bm{x});G)\;,$
$g^{m}_{\theta}(\bm{x};G)=\max_{k\in\{1,\dots,K\}}\text{m\_softmax}_{k}(f_{\theta}(\bm{x});G)\;,$
where the functions $h^{m}_{\theta}$ and $g^{m}_{\theta}$ correspond to $h_{\theta}$ and $g_{\theta}$ but restricting the choice of viable clusters to $\{1,\dots,K\}\setminus G$. In particular, the function $h^{m}_{\theta}$ computes the cluster assignment for a datapoint $\bm{x}$ restricting to clusters in $\{1,\dots,K\}\setminus G$, and $g^{m}_{\theta}$ computes the corresponding predicted probability. Lastly, we define
$\mathcal{D}^{c}:=\{\bm{x}\in\mathcal{D}\mid h_{\theta}(\bm{x})=c\}\;,$
i.e., the subset of data points assigned to a given cluster $c\in\{1,\dots,K\}$. Similarly, we denote as $\mathcal{D}^{G}=\cup_{c\in G}D^{c}$ the subset of data points assigned to a group of clusters $G\subset\{1,\dots,K\}$.
Algorithm 1 Logits to Hierarchies (L2H). Given aggregation function $\Lambda$, and functions $g_{\theta},g^{m}_{\theta}$ defined as above for pre-trained $K$-clustering model $f_{\theta}$.
  Input: Dataset $\mathcal{D}$.
@@
  # Groups initialized as single clusters
  Initialize groups $\mathcal{G}=\{G_{1},\dots,G_{K}\}$ where $G_{k}=\{k\}$
  # Hierarchy initialized as empty list
  H = []
  for step $t$ from 11 to K1K-1 do
@@
  # Select group with lowest score for merging
  Take $G^{\star}\in\text{argmin}_{G\in\mathcal{G}}s(G)$
@@
    # Compute total pred. probability per cluster as in Eqn.2
   Compute $rp(c):=\sum_{\begin{subarray}{c}\bm{x}\in\mathcal{D}^{G^{*}}\\ h_{\theta}^{m}(\bm{x};G^{*})=c\end{subarray}}\!\!\!\!g_{\theta}^{m}(\bm{x};G^{*})$
@@
  # Select $G^{\dagger}$, most related group to $G^{*}$, for merging
  Take $G^{\dagger}=\text{argmax}_{G\in\mathcal{G}\setminus\{G^{*}\}}\frac{1}{|G|}\sum_{c\in G}rp(c)$
  # Update groups
  Update $\mathcal{G}$ by merging groups $G^{*}$ and $G^{\dagger}$
  # Update hierarchy
  Update H by adding that $G^{*}$ and $G^{\dagger}$ are merged at step $t$
@@
We describe our proposed method in Algorithm 1. At the start of the procedure, $K$ groups are initialized as single clusters.  [^2] Note that here cluster is used to refer to a single cluster found by the pre-trained model, while group refers to a set of clusters that are grouped together at a given step of the algorithm. At each iteration, two groups are merged into a single group, constructing a tree of clusters up to the root in K1K-1 iterations. Each iteration can be split into two stages. In the first stage, a score is computed for each group. To compute the score for a given group $G$ we aggregate the predicted probabilities for the data points assigned to clusters contained in $G$ as
@@
where $\Lambda$ is a chosen aggregation function (e.g., the sum). Then, the lowest-scored group $G^{*}$ is selected for merging at this iteration, which concludes the first stage.
In the second stage, we search for the group $G^{\dagger}$ that is most related to $G^{*}$ to perform the merging. To do so, as mentioned above, we look at the subset of data points assigned to clusters in $G^{*}$: for these data points, we recompute cluster assignments and predicted probabilities, this time restricting to clusters not contained in $G^{*}$. More formally, the total reassigned predicted probability to each cluster not contained in $G^{*}$ is computed as
$rp(c):=\sum_{\begin{subarray}{c}\bm{x}\in\mathcal{D}^{G^{*}}\\ h_{\theta}^{m}(\bm{x};G^{*})=c\end{subarray}}g_{\theta}^{m}(\bm{x};G^{*})\;,\quad\forall c\in\{1,..,K\}\setminus G^{*}\;.$ (2)
Note that this quantity can be interpreted as a measure of relatedness between each cluster $c\in\{1,..,K\}\setminus G^{*}$, and the group of clusters $G^{*}$. The most related group to $G^{*}$ is finally selected as $G^{\dagger}\in\text{argmax}_{G\in\mathcal{G}\setminus\{G^{*}\}}\frac{1}{|G|}\sum_{c\in G}rp(c)$, i.e., by averaging the total reassigned predicted probability across clusters in each group and selecting the group with the highest average.
Given that cluster assignments and corresponding predicted probabilities can be computed via simple operations on the logits, the whole procedure can be executed given only the logits from a pre-trained model $f_{\theta}$ for the dataset $\mathcal{D}$. For further clarification, we provide a graphical illustration of the proposed grouping strategy (Figure 1) and an example Python implementation in Appendix A.